OpenStudy (anonymous):

Solve the system of linear equations below. x = -12, y = 7 x = 3, y = 2 x = 6, y = 1 x = 6, y = 2

OpenStudy (anonymous):

The Linear Equations Are x − 3y = -3 x + 3y = 9

OpenStudy (anonymous):

The Linear Equations Are x − 3y = -3 x + 3y = 9

OpenStudy (anonymous):